An investment will pay $50 at the end of each of the next 3 years, $250 at the end of Year 4, $400 at the end of Year 5, and $550 at the end of Year 6. If other investments of equal risk earn 9% annually, what is its present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present value: $

Future value:
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=50/1.09+50/1.09^2+50/1.09^3+250/1.09^4+400/1.09^5+550/1.09^6

=$891.59(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value=891.59*(1.09)^6

=$1495.29(Approx).
Answer #2

SOLUTION :


Discount rate, r = 09% = 0.09 

=> 1 + r = 1.09


PV 

= 50/1.09+50/1.09^2+50/1.09^3+250/1.09^4+400/1.09^5+550/1.09^6

= 891.59 ($) (ANSWER).


FV 

= PV*1.09^6 

= 891*1.09^6

= 1495.29 ($) (ANSWER)

answered by: Tulsiram Garg